# Gradient Descent on Logistic Regression with Non-Separable Data and Large Step Sizes

### v1

@@ 1 Introduction
where $\eta_{t}$ is the step size. As ${\mathcal{L}}$ is convex and $L$-smooth, classical optimization theory guarantees that a constant step size $\eta_{t}=\eta<2/L$ is sufficient for GD to converge to $w^{*}$ (Nesterov, 2018).
@@
If the data is not linearly-separable, the objective is strictly convex as long as the features have full-rank, thus the unique minimizer $w^{*}$ is finite. For this reason, one can not expect convergence under an arbitrarily large step size: indeed, classical dynamical systems theory (Strogatz, 2018; Sayama, 2015) shows that $w^{*}$ becomes unstable when the step size $\eta>2/\lambda$, where $\lambda$ is the largest eigenvalue of the Hessian of $\mathcal{L}$ at $w^{*}$. A natural question to ask is whether this is the only barrier in the non-separable case: can we still guarantee convergence, as in the separable setting, using a “large” step size $\eta$ between $2/L$ and the “critical” step size of $2/\lambda$? And what happens at even larger $\eta$? Large step sizes are interesting because they can often lead to faster convergence, both for logistic regression in the separable case (Axiotis and Sviridenko, 2023; Wu et al., 2024) and for more general problems (Altschuler and Parrilo, 2023; Grimmer et al., 2023; Mishkin et al., 2024; Oymak, 2021; Wu and Su, 2023). We also don’t know a priori whether the data is separable, so it would be helpful to gain a better understanding of what happens when we push the step size beyond the $2/L$ limit. The large step size regime has also been studied for deep neural networks, often referred to as the Edge-of-Stability (Cohen et al., 2021), and is known to cause spikes or catapults in the initial steps of optimization (Zhu et al., 2023a).
@@ 2 Background
Non-separable logistic regression problems differ from the separable setting largely due to the location of the minimizer. If the data is non-separable, the objective is strictly convex in the subspace spanned by the features $\{x_{i}\}$, and the solution is no longer attained at infinity. GD on logistic regression is essentially a discrete time nonlinear dynamical system for which $w^{*}$ is a fixed point. A necessary condition for $w^{*}$ to be locally (linearly) stable is to have a step size smaller than $2/\lambda$, where $\lambda$ is the largest eigenvalue of the Hessian at $w^{*}$ (Strogatz, 2018; Sayama, 2015). Local stability means GD converges to $w^{*}$ when we initialize close enough to it. On the other hand, if we view the problem from a convex optimization perspective, a sufficient condition for global convergence is to require $\eta<2/L$, where $L$ is a global upper bound on the Hessian, which can be much larger than $\lambda$. One can relax the $2/L$ requirement by using step sizes that depend on the local smoothness (Hessian around the current iterate), as do Ji and Telgarsky (2019) for non-separable logistic regression and Mishkin et al. (2024) for general convex problems. However, these step sizes still effectively requires that the objective decreases monotonically, which is not guaranteed with a constant step size greater than $2/L$. For logistic regression, Liu et al. (2023) created a two-example dataset with identical features but opposite labels, on which GD can enter a stable period-$2$ cycle when the step size is greater than a critical value. Unfortunately, this critical step size coincides with both $2/L$ and $2/\lambda$ due to the degeneracy of the dataset. which means there is still little known about what happens when there is a non-trivial gap between $2/L$ and $2/\lambda$.
@@
Beyond linear classification problems, period-doubling bifurcations and chaos in GD dynamics under large step sizes have been observed in many problem settings. For least squares problems, van den Doel and Ascher (2012) showed that occasionally taking very large step sizes can lead to much faster convergence. But if these step sizes are too large, GD can behave chaotically. Beyond linear models, Chen and Bruna (2022) gave sufficient conditions for a period-$2$ cycle to exist for one-dimensional loss functions, but the study is mostly restricted to the squared loss. Zhu et al. (2022, 2023a) empirically studied non-monotonic convergence of GD under the critical step size $\eta<\lambda_{0}$, where $\lambda_{0}$ is the largest eigenvalue of the Hessian at initialization. Their studies apply to neural networks in the NTK regime but are also limited to the squared loss. Chen et al. (2023) proved that under restrictive conditions on the input data and architecture, GD on neural networks with nonlinear activations boils down to a one-dimensional cubic map that can behave periodically or chaotically. Once again, these results only apply to the squared loss. Under general loss functions and model architectures, Ahn et al. (2022b) gave intuitions to when GD can converge under unstably large step sizes, while Danovski et al. (2024) observed stable oscillation and chaos for neural network training.
@@
In the deep learning literature, convergence of GD under large step sizes is commonly known as the Edge of Stability (EoS) phenomenon (Cohen et al., 2021). Specifically, it has been observed that when training neural networks, the largest eigenvalue of the Hessian, also referred to as the sharpness, often hovers right at, or even above $2/\eta$. Zhu et al. (2023b) illustrated this phenomenon using a minimalist 4-parameter scalar network with the quadratic loss, where GD iterates initially oscillate, then de-bifurcate, leading to convergence at an EoS minimum. Another motivation for studying GD step sizes in the EoS regime is that large step sizes can be crucial in learning the underlying representations of the problem. For instance, Ahn et al. (2022a) showed that in a sparse-coding setup, one can only learn the bias term necessary for recovery by dialing up the step size into the unstable regime. While there are many more works studying the EoS regime for non-convex problems (Song and Yun, 2023; Kreisler et al., 2023; Wang et al., 2023), we believe it is useful to take a step back and closely examine what exactly happens on just linear models, especially for the logistic loss which seems to be under-explored. We show that stable cycles can occur under the critical step size $2/\lambda$, and illustrate precisely how these cycles arise.
@@ 3 Period-doubling bifurcation and chaos
When the examples are not linearly-separable, that is, for all $w\in\mathbb{R}^{d}$, there exists $i\in\{1,\dots,n\}$ such that $y_{i}w^{\mathsf{\scriptscriptstyle T}}x_{i}<0$, the logistic regression objective ${\mathcal{L}}$ in Equation 1 is strictly convex, as long as the $x_{i}$’s have full rank. The solution $w^{*}$ is necessarily unique and finite, so we can simply run GD with increasing step sizes to examine its convergence properties. In Figure 1, we see that GD is convergent for small step sizes, up to a point at which a period-$2$ cycle emerges. As we continue to increase the step size beyond this point, a sequence of period-doubling bifurcation occurs, and GD converges to cycles of longer periods. Eventually, this period-doubling cascades into chaos.
Figure 1: Bifurcation diagrams on two binary classification datasets from Chang and Lin (2011), both are non-separable. For each step size, we run GD for $T=5\cdot 10^{5}$ iterations with $1024$ different random initializations of varying scales. Each point corresponds to the loss (first row) or the (scaled) largest eigenvalue of the Hessian (second row) evaluated at the final iterate $w_{T}$. When multiple points are visible, GD either converged to a cycle or is chaotic under that step size.
@@
The fact that GD undergoes period-doubling bifurcation on the logistic regression objective is perhaps not so surprising — this phenomenon has long been observed and studied on numerous nonlinear maps (Strogatz, 2018). Unlike GD on linear regression which results in a linear map in discrete time, the sigmoid link in the gradient adds a nonlinearity that gives rise to a much richer spectrum of behaviour. What the bifurcation diagrams can help us see is at what step size we cease to have convergence to $w^{*}$. To further our understanding, consider the gradient and Hessian of 1:
$\nabla{\mathcal{L}}(w)=\frac{1}{n}\sum_{i=1}^{n}\sigma(-w^{\mathsf{\scriptscriptstyle T}}x_{i})(-x_{i}),\quad\nabla^{2}{\mathcal{L}}(w)=\frac{1}{n}\sum_{i=1}^{n}\sigma^{\prime}(-w^{\mathsf{\scriptscriptstyle T}}x_{i})x_{i}x_{i}^{\mathsf{\scriptscriptstyle T}},$ (3)
where $\sigma(z)=1/(1+\exp(-z))$ is the sigmoid function. We will use $\lambda$ to denote $\lambda_{\max}(\nabla^{2}{\mathcal{L}}(w^{*}))$. As the Hessian is maximized at $w=0$, the global upper bound on the Hessian is given by $L=XX^{\mathsf{\scriptscriptstyle T}}/(4n)$, where $X\in\mathbb{R}^{n\times d}$ is the feature matrix. One interesting observation from Figure 1 lies in the second row, where we plot the final values of $\lambda_{\max}(\nabla^{2}{\mathcal{L}}(w_{T}))$, scaled by $\eta/2$. For all step sizes smaller than $2/L$, this value is clearly convergent, as $w_{t}$ converges to $w^{*}$ regardless of initialization. It appears that GD remains convergent beyond $\eta=2/L$ until $\eta=2/\lambda$, at which period-$2$ cycles begin to appear. This is reasonable as $w^{*}$ is the fixed point of the GD map $T(w)=w-\eta\nabla{\mathcal{L}}(w)$, and first-order stability of $w^{*}$ is guaranteed if the eigenvalues of the Jacobian of $T$
@@
lie strictly within the unit circle (Sayama, 2015, Chapter 5.7). As a result, $\eta<2/\lambda$ is a necessary condition for GD to converge to $w^{*}$. Next, we illustrate on a toy dataset that the gap between $2/L$ and $2/\lambda$ can grow arbitrarily large, and that analyzing the cycles is very challenging.
@@ 3.1 A toy dataset
Consider $n\geq 2$ examples such that $y_{i}=1$ for all $i$. Let $v$ be an arbitrary point on the $d$-dimensional unit sphere. The dataset consists of $n-1$ copies of $v$, and a single copy of $-v$. Clearly, this dataset is not separable by any linear classifier that goes through the origin. The gradient and Hessian simplifies to (see Appendix C)
@@
while globally $L=1/4$. So for large $n$, the gap between $2/L$ and $2/\lambda$ grows quickly.
@@
For $n=2$ and $\eta\geq 2/\lambda=8$, the two points of oscillation are given by the two values of
@@ 4 Technical setup
where $y_{i}\in\left\{-1,1\right\}$, $x_{i}\in\mathbb{R}^{d}$, and $\ell$ is the loss function on a single example. We are particularly interested in the case where the data is not linearly-separable.
Definition 1 (Non-separability).
For all $w\in\mathbb{R}^{d}$, there exists $i\in\left\{1,\dots,n\right\}$ such that $y_{i}w^{\mathsf{\scriptscriptstyle T}}x_{i}<0$.
While our motivation is to study the behaviour of GD on logistic regression under large step sizes, our results will be stated in terms of general loss functions $\ell$ satisfying the following set of assumptions.
@@
The loss function on the individual examples $\ell:\mathbb{R}\to\mathbb{R}_{+}$ satisfies:
@@
    $\ell$ is three-times continuously-differentiable and strictly convex, and $\lim_{z\to-\infty}\ell(z)=0$.
@@
    $\lim_{z\to-\infty}\ell^{\prime}(z)=0$ and $\lim_{z\to\infty}\ell^{\prime}(z)=1$,
@@
    $\ell^{\prime\prime}$ is increasing on $(-\infty,0]$ and decreasing on $[0,\infty)$; furthermore, it decays fast enough that
@@
The logistic loss $\ell(z)=\log(1+\exp(z))$ and the squareplus function $\ell(z)=0.5(\sqrt{4+z^{2}}+z)$ (Barron, 2021) both satisfy Assumption 1, as verified in Section D.1. Outside of these two losses, there are few commonly-used machine learning loss functions that satisfy our assumptions. The purpose of stating our results in this way is to emphasize that our results are not an algebraic consequence of the logistic loss, but rather, a consequence of its structural properties.
@@ 5 One-dimensional case
As discussed in the previous section, although $\eta<2/\lambda$ is a necessary condition for convergence, it is not sufficient. Specifically, it only guarantees local convergence to $w^{*}$ when we initialize within a neighborhood of $w^{*}$. From Figure 1, it appears that we do converge globally for all $\eta<2/\lambda$, as the initializations we used vary in scale. However, as we show later, this is not the case for all datasets. The first natural question to ask is what is a sufficient condition on the step size to guarantee global convergence? As it turns out, when $d=1$, this step size is given by $\eta<1/\lambda$.
@@
Suppose Assumption 1 holds for the classification problem in $1$ dimension with non-separable data, that is, $w^{*}$ is finite. Then the GD iterates $\left\{w_{t}\right\}$ converge to $w^{*}$ when $\eta\leq 1/{\mathcal{L}}^{\prime\prime}(w^{*})$ for any initialization $w_{0}$. Moreover, if we choose $\eta=1/{\mathcal{L}}^{\prime\prime}(w^{*})$, then for all $\epsilon>0$, it holds that $(w_{\tau}-w^{*})^{2}\leq\epsilon$ for some
$\tau\leq 1+\max\left\{0,-w_{0}\cdot\operatorname{sign}(w^{*})\cdot{\mathcal{L}}^{\prime\prime}(w^{*})/{\mathcal{L}}^{\prime\prime}(0)\right\}+{\mathcal{O}}(\log(1/\epsilon)).$
Convergence is straightforward. Suppose $w^{*}>0$. The interval ${\mathcal{I}}=[w^{*},+\infty)$ is an invariant set under the step size requirement, that is, if $w_{t}\in{\mathcal{I}}$, then $w_{t+1}\in{\mathcal{I}}$. Furthermore, $w_{t}$ converges to $w^{*}$ on ${\mathcal{I}}$. So either we converge directly in ${\mathcal{I}}$, or we initialize in $(-\infty,w^{*})$, from which we either approach $w^{*}$ form the left, or cross over into ${\mathcal{I}}$ , within which convergence is guaranteed. The detailed proof including the rate can be found in Section A.1. The next result shows that if we increase the step size beyond $1/{\mathcal{L}}^{\prime\prime}(w^{*})$, global convergence is no longer guaranteed.
@@
Suppose Assumption 1 holds and $d=1$. Then for all $\gamma\in(1,2]$, there exists a non-separable classification problem on which a GD trajectory under the step size $\eta=\gamma/{\mathcal{L}}^{\prime\prime}(w^{*})$ converges to a cycle of period $k>1$.
It is worth emphasizing that Theorem 2 does not imply that a cycle is possible for every dataset. It only shows that if we were to first pick a $\gamma\in(1,2]$, then we can construct a dataset such that GD with step size $\eta=\gamma/{\mathcal{L}}^{\prime\prime}(w^{*})$ can converge to a cycle. Note that this construction implies $\eta>2/L$ where $L={\mathcal{L}}^{\prime\prime}(0)$ is the global smoothness constant, as otherwise we would have global convergence. Stability of the cycle coexists with stability of the solution $w^{*}$. Therefore, depending on the initialization, GD may still converge to $w^{*}$ with the same step size. Moreover, given an arbitrary dataset, it is not always easy to verify whether a cycle exists before hitting the $2/\lambda$ step size. In fact, when we ran GD using many different initializations of varying scales in on real datasets (Figure 1), we did not see a clear cycle emerging at all until $\eta=2/\lambda$. We now present the main proof steps.
@@
$w_{k}<0<w_{k-1}<\dots<w_{0}$ (9)
exists. Each iterate is generated via the GD update on ${\mathcal{L}}$ with step size $\eta=\gamma/{\mathcal{L}}^{\prime\prime}(w^{*})$, illustrated in the left panel of Figure 4. Furthermore, we can ensure that $0<w_{k+1}<w_{0}$.
@@
${\mathcal{L}}_{\text{perturbed}}(w)={\mathcal{L}}(w)+\frac{w_{0}-w_{k+1}}{\eta}\max\left\{-w,\,0\right\}.$ (10)
Observe that the ReLU does not contribute any additional curvature on top of ${\mathcal{L}}$, and it also does not contribute any gradient to all points $w>0$. Therefore, the minimizer remains the same. So if we were to run GD starting at the same $w_{0}$ on ${\mathcal{L}}_{\text{perturbed}}$ with the step size $\eta_{\text{perturbed}}=\gamma/{\mathcal{L}}^{\prime\prime}_{\text{perturbed}}(w^{*})=\eta$, the new trajectory would coincide with the original trajectory up to step $k$. Let $\tilde{w}_{t}$ denote the iterates of this new trajectory for $t\geq 0$. At point $\tilde{w}_{k}=w_{k}<0$, the ReLU becomes active. Applying one GD update at this point leads to
$\tilde{w}_{k+1}$ $=\tilde{w}_{k}-\eta{\mathcal{L}}^{\prime}_{\text{perturbed}}(\tilde{w}_{k})$ (11)
$=\tilde{w}_{k}-\eta{\mathcal{L}}^{\prime}(\tilde{w}_{k})+\eta\frac{w_{0}-w_{k+1}}{\eta}$
$=w_{k+1}+w_{0}-w_{k+1}$

### v2

@@ 1 Introduction
where $\eta_{t}$ is the step size. As ${\mathcal{L}}$ is convex and $L$-smooth, classical optimization theory guarantees that a constant step size $\eta_{t}=\eta<2/L$ is sufficient for GD to converge to $w^{*}$ for any initialization (Nesterov, 2018).
@@
If the data is not linearly-separable, the objective is strictly convex as long as the features have full-rank, thus the unique minimizer $w^{*}$ is finite. For this reason, one can not expect convergence under an arbitrarily large step size: indeed, classical dynamical systems theory (Strogatz, 2018; Sayama, 2015) shows that $w^{*}$ becomes unstable when the step size $\eta>2/\lambda$, where $\lambda$ is the largest eigenvalue of the Hessian of $\mathcal{L}$ at $w^{*}$. A natural question to ask is whether this is the only barrier in the non-separable case: can we still guarantee convergence, as in the separable setting, for all “large” step sizes $\eta$ between $2/L$ and the “critical” step size of $2/\lambda$? And what happens at even larger $\eta$? Large step sizes are interesting because they can often lead to faster convergence, both for logistic regression in the separable case (Axiotis and Sviridenko, 2023; Wu et al., 2024) and for more general problems (Altschuler and Parrilo, 2023; Grimmer et al., 2023; Mishkin et al., 2024; Oymak, 2021; Wu and Su, 2023; Ahn et al., 2022). We also don’t know a priori whether the data is separable, so it would be helpful to gain a better understanding of what happens when we push the step size beyond the $2/L$ limit. The large step size regime has also been studied for deep neural networks, often referred to as the Edge-of-Stability (Cohen et al., 2021), and is known to cause spikes or catapults in the initial steps of optimization (Zhu et al., 2024).
@@ 2 Background
Non-separable logistic regression problems differ from the separable setting largely due to the location of the minimizer. If the data is non-separable, the objective is strictly convex in the subspace spanned by the features $\{x_{i}\}$, and the solution is no longer attained at infinity. GD on logistic regression is essentially a discrete time nonlinear dynamical system for which $w^{*}$ is a fixed point. A necessary condition for $w^{*}$ to be locally (linearly) stable is to have a step size smaller than $2/\lambda$, where $\lambda$ is the largest eigenvalue of the Hessian at $w^{*}$ (Strogatz, 2018; Sayama, 2015). Local stability means GD converges to $w^{*}$ when we initialize close enough to it. On the other hand, if we view the problem from a convex optimization perspective, a sufficient condition for global convergence is to require $\eta<2/L$, where $L$ is a global upper bound on the Hessian, which can be much larger than $\lambda$. One can relax the $2/L$ requirement by leveraging the generalized self-concordance property of logistic regression (Bach, 2009), yielding an instance-dependent large step size. One can also use step sizes that depend on the local smoothness (Hessian around the current iterate), as do Ji and Telgarsky (2019) for non-separable logistic regression and Mishkin et al. (2024) for general convex problems. However, these step sizes are either still strictly below the $2/\lambda$ stability threshold, or still effectively require that the objective decreases monotonically, which is not guaranteed for $\eta>2/L$. For logistic regression, Liu et al. (2023) created a two-example dataset with identical features but opposite labels, on which GD can enter a stable period-$2$ cycle when the step size is greater than a critical value. Unfortunately, this critical step size coincides with both $2/L$ and $2/\lambda$ due to the degeneracy of the dataset. which means there is still little known about what happens when there is a non-trivial gap between $2/L$ and $2/\lambda$.
@@
Beyond linear classification problems, period-doubling bifurcations and chaos in GD dynamics under large step sizes have been observed in many problem settings. For least squares problems, van den Doel and Ascher (2012) showed that occasionally taking very large step sizes can lead to much faster convergence. But if these step sizes are too large, GD can behave chaotically. Beyond linear models, Chen and Bruna (2022) gave sufficient conditions for a period-$2$ cycle to exist for one-dimensional loss functions, but the study is mostly restricted to the squared loss. Zhu et al. (2022, 2024) empirically studied non-monotonic convergence of GD under the critical step size $\eta<1/\lambda_{0}$, where $\lambda_{0}$ is the largest eigenvalue of the Hessian at initialization.Their studies apply to general neural networks but are also limited to the squared loss. Chen et al. (2024) proved that under restrictive conditions on the input data and architecture, GD on neural networks with nonlinear activations boils down to a one-dimensional cubic map that can behave periodically or chaotically. Once again, these results only apply to the squared loss. Under general loss functions and model architectures, Ahn et al. (2022) gave intuitions to when GD can converge under unstably large step sizes, while Danovski et al. (2024) observed stable oscillation and chaos for neural network training.
@@
In the deep learning literature, convergence of GD under large step sizes is commonly known as the Edge of Stability (EoS) phenomenon (Cohen et al., 2021). Specifically, it has been observed that when training neural networks, the largest eigenvalue of the Hessian, also referred to as the sharpness, often hovers right at, or even above $2/\eta$, while the objective continues to decrease. Zhu et al. (2023) illustrated this phenomenon using a minimalist 4-parameter scalar network with the quadratic loss, where GD iterates initially oscillate, then de-bifurcate, leading to convergence at an EoS minimum. Another motivation for studying GD step sizes in the EoS regime is that large step sizes can be crucial in learning the underlying representations of the problem. For instance, Ahn et al. (2023) showed that in a sparse-coding setup, one can only learn the bias term necessary for recovery by dialing up the step size into the unstable regime. While there are many more works studying the EoS regime for non-convex problems (Song and Yun, 2023; Kreisler et al., 2023; Wang et al., 2023; Lu et al., 2024), we believe it is useful to take a step back and closely examine what exactly happens on just linear models, especially for the logistic loss which seems to be under-explored. We show that stable cycles can occur under the critical step size $2/\lambda$, and illustrate precisely how these cycles arise.
@@ 3 Period-doubling bifurcation and chaos
When the examples are not linearly-separable, that is, for all $w\in\mathbb{R}^{d}$, there exists $i\in\{1,\dots,n\}$ such that $y_{i}w^{\mathsf{\scriptscriptstyle T}}x_{i}\leq 0$, the logistic regression objective ${\mathcal{L}}$ in Equation 1 is strictly convex, as long as the $x_{i}$’s have full rank. The solution $w^{*}$ is necessarily unique and finite, so we can simply run GD with increasing step sizes to examine its convergence properties. In Figure 1, we see that GD is convergent for small step sizes, up to a point at which a period-$2$ cycle emerges. As we continue to increase the step size beyond this point, a sequence of period-doubling bifurcation occurs, and GD converges to cycles of longer periods. Eventually, this period-doubling cascades into chaos.
Figure 1: Bifurcation diagrams on two binary classification datasets from the LIBSVM repository (Chang and Lin, 2011), both are non-separable. For each step size, we run GD for $T=5\cdot 10^{5}$ iterations with $1024$ different random initializations of varying scales. Each point corresponds to the loss (first row) or the (scaled) largest eigenvalue of the Hessian (second row) evaluated at the final iterate $w_{T}$. When multiple points are visible, GD either converged to a cycle or is chaotic under that step size.
@@
The fact that GD undergoes period-doubling bifurcation on the logistic regression objective is perhaps not so surprising — this phenomenon has long been observed and studied on numerous nonlinear maps (Strogatz, 2018). Unlike GD on linear regression which results in a linear map in discrete time, the sigmoid link in the gradient adds a nonlinearity that gives rise to a much richer spectrum of behaviour. What the bifurcation diagrams can help us see is at what step size we cease to have convergence to $w^{*}$. To further our understanding, consider the gradient and Hessian of Equation 1:
$\nabla{\mathcal{L}}(w)=\frac{1}{n}\sum_{i=1}^{n}\sigma(-y_{i}w^{\mathsf{\scriptscriptstyle T}}x_{i})(-y_{i}x_{i}),\quad\nabla^{2}{\mathcal{L}}(w)=\frac{1}{n}\sum_{i=1}^{n}\sigma^{\prime}(-y_{i}w^{\mathsf{\scriptscriptstyle T}}x_{i})x_{i}x_{i}^{\mathsf{\scriptscriptstyle T}},$ (3)
where $\sigma(z)=1/(1+\exp(-z))$ is the sigmoid function. We will use $\lambda$ to denote $\lambda_{\max}(\nabla^{2}{\mathcal{L}}(w^{*}))$. As the Hessian is maximized at $w=0$, the global upper bound on the Hessian is given by $L=\lambda_{\max}(XX^{\mathsf{\scriptscriptstyle T}}/(4n))$, where $X\in\mathbb{R}^{n\times d}$ is the feature matrix. One interesting observation from Figure 1 lies in the second row, where we plot the final values of $\lambda_{\max}(\nabla^{2}{\mathcal{L}}(w_{T}))$ scaled by $\eta/2$. For all step sizes below $2/L$, this value is clearly convergent, as $w_{t}$ converges to $w^{*}$ regardless of initialization. It appears that GD remains convergent beyond $\eta=2/L$ until $\eta=2/\lambda$, at which period-$2$ cycles begin to appear. This is reasonable as $w^{*}$ is the fixed point of the GD map $T(w)=w-\eta\nabla{\mathcal{L}}(w)$, and first-order stability of $w^{*}$ is guaranteed if the eigenvalues of the Jacobian of $T$
@@
lie strictly within the unit circle (Sayama, 2015, Chapter 5.7). As a result, $\eta<2/\lambda$ is a necessary condition for GD to converge to $w^{*}$ in a neighborhood of $w^{*}$. Next, we illustrate on a toy dataset that the gap between $2/L$ and $2/\lambda$ can grow arbitrarily large, and that analyzing the cycles is very challenging.
@@ 3.1 A toy dataset
Consider $n\geq 2$ examples such that $y_{i}=1$ for all $i$. Let $v$ be an arbitrary point on the $d$-dimensional unit sphere. The dataset consists of $n-1$ copies of $v$, and a single copy of $-v$. Clearly, this dataset is not separable by any linear classifier that goes through the origin. The gradient and Hessian simplify to (see Appendix D)
@@
while $L=1/4$. So for large $n$, the gap between the step sizes $2/L$ and $2/\lambda$ grows quickly.
@@
The bifurcation diagrams of GD on this toy dataset are presented in Figure 8. For $n=2$ and $\eta\geq 2/\lambda=8$, the two points of oscillation are given by the two values of
@@ 4 Technical setup
where $y_{i}\in\left\{-1,1\right\}$, $x_{i}\in\mathbb{R}^{d}$, and $\ell$ is the loss function on a single example. We use $\lambda$ to denote $\lambda_{\max}(\nabla^{2}{\mathcal{L}}(w^{*}))$. We are particularly interested in the case where the data is not linearly-separable. Concretely, we assume that for all $w\in\mathbb{R}^{d}$, there exists $i\in\left\{1,\dots,n\right\}$ such that $y_{i}w^{\mathsf{\scriptscriptstyle T}}x_{i}\leq 0$. This implies that $\nabla^{2}{\mathcal{L}}(w)\neq 0$ for all $w$. While our motivation is to study the behaviour of large step size GD on logistic regression, our results will be stated in terms of general loss functions $\ell$ satisfying the following set of assumptions.
@@
The individual loss $\ell:\mathbb{R}\to\mathbb{R}_{+}$ is
@@
    three-times continuously-differentiable and strictly convex, and $\lim_{z\to-\infty}\ell(z)=0$,
@@
    $\lim_{z\to-\infty}\ell^{\prime}(z)=0$ and $\lim_{z\to\infty}\ell^{\prime}(z)=1$, and
@@
    $\ell^{\prime\prime}$ is increasing on $(-\infty,0]$ and decreasing on $[0,\infty)$; furthermore, it decays rapidly as
@@
The logistic loss $\ell(z)=\log(1+\exp(z))$ and the squareplus function $\ell(z)=0.5(\sqrt{4+z^{2}}+z)$ (Barron, 2021) both satisfy Assumption 1, as verified in C.1. Outside of these two losses, there are few commonly-used machine learning loss functions that satisfy our assumptions. The purpose of stating our results in this way is to emphasize that our results are not an algebraic consequence of the logistic loss, but rather, a consequence of its structural properties.
@@ 5 One-dimensional case
As discussed in the previous section, although $\eta<2/\lambda$ is a necessary condition for convergence, it is not sufficient. Specifically, it only guarantees local convergence to $w^{*}$ when we initialize within a neighborhood of $w^{*}$. From Figure 1, it appears that we do converge globally for all $\eta<2/\lambda$, as the initializations we used vary in scale. However, as we show later, this is not the case for all datasets. The first natural question to ask is what is a sufficient condition on the step size to guarantee global convergence? As it turns out, when $d=1$, this step size is given by $\eta\leq 1/\lambda$.
@@
Suppose Assumption 1 holds for the classification problem in $1$ dimension with non-separable data. Then the GD iterates $\left\{w_{t}\right\}$ under $\eta\leq 1/{\mathcal{L}}^{\prime\prime}(w^{*})$ converge to $w^{*}$ for any initialization $w_{0}$. When $\eta=1/{\mathcal{L}}^{\prime\prime}(w^{*})$, let $\bar{\tau}\coloneqq 1+\max\left\{0,-\lceil w_{0}\cdot\operatorname{sign}(w^{*})\cdot{\mathcal{L}}^{\prime\prime}(w^{*})/{\mathcal{L}}^{\prime\prime}(0)\rceil\right\}$. Then there exists $\rho\in(0,1)$ such that for all $t\geq\bar{\tau}$, we have $(w_{t+1}-w^{*})^{2}\leq\rho(w_{t}-w^{*})^{2}$.
Convergence is straightforward. Suppose $w^{*}>0$. The interval ${\mathcal{I}}=[w^{*},+\infty)$ is an invariant set under the step size requirement, that is, if $w_{t}\in{\mathcal{I}}$, then $w_{t+1}\in{\mathcal{I}}$. Furthermore, $w_{t}$ converges to $w^{*}$ on ${\mathcal{I}}$. So either we converge directly in ${\mathcal{I}}$, or we initialize in $(-\infty,w^{*})$, from which we either approach $w^{*}$ from the left, or cross over into ${\mathcal{I}}$ , within which convergence is guaranteed. The detailed proof including the rate can be found in Section A.1. The next result shows that if we increase the step size beyond $1/{\mathcal{L}}^{\prime\prime}(w^{*})$, global convergence is no longer guaranteed.
@@
For any $\ell$ for which Assumption 1 holds and any $\gamma\in(1,2]$, there exists a 1-dimensional non-separable classification problem of the form 7, on which there exists a GD trajectory under the step size $\eta=\gamma/{\mathcal{L}}^{\prime\prime}(w^{*})$ that converges to a cycle of period $k>1$.
It is worth emphasizing that Theorem 2 does not imply a cycle is possible for every dataset. It only shows that if we were to first pick a $\gamma\in(1,2]$, then we can construct a dataset such that GD with step size $\eta=\gamma/{\mathcal{L}}^{\prime\prime}(w^{*})$ can converge to a cycle. Note that this construction necessarily implies $\eta>2/L$ where $L={\mathcal{L}}^{\prime\prime}(0)$ is the global smoothness constant, as otherwise we would have global convergence. Stability of the cycle coexists with stability of the solution $w^{*}$. Therefore, depending on the initialization, GD may still converge to $w^{*}$ with the same step size. Moreover, given an arbitrary dataset, it is not always easy to verify whether a cycle exists before hitting the $2/\lambda$ step size. In fact, when we ran GD using many different initializations of varying scales in on real datasets (Figure 1), we did not see a clear cycle emerging at all until the stability threshold of $\eta=2/\lambda$. We now state the main proof steps.
@@
$w_{k-1}<0<w^{*}<w_{k-2}<\dots<w_{0}$ (9)
exists. Each iterate is generated via the GD update on ${\mathcal{L}}$ with step size $\eta=\gamma/{\mathcal{L}}^{\prime\prime}(w^{*})$, illustrated in the left panel of Figure 4. Furthermore, we can ensure that $w^{*}<w_{k}<w_{0}$.
@@
${\mathcal{L}}_{\text{perturbed}}(w)={\mathcal{L}}(w)+\frac{w_{0}-w_{k}}{\eta}\max\left\{-w,\,0\right\}.$ (10)
Observe that the ReLU does not contribute any additional curvature on top of ${\mathcal{L}}$, and it also does not contribute any gradient to all points $w>0$. Therefore, the minimizer remains the same. So if we were to run GD starting at the same $w_{0}$ on ${\mathcal{L}}_{\text{perturbed}}$ with the step size $\eta_{\text{perturbed}}=\gamma/{\mathcal{L}}^{\prime\prime}_{\text{perturbed}}(w^{*})=\eta$, the new trajectory would coincide with the original trajectory up to step $k$. Let $\tilde{w}_{t}$ denote the iterates of this new trajectory for $t\geq 0$. At point $\tilde{w}_{k-1}=w_{k-1}<0$, the ReLU becomes active. Applying one GD update at this point leads to
$\tilde{w}_{k}$ $=\tilde{w}_{k-1}-\eta{\mathcal{L}}^{\prime}_{\text{perturbed}}(\tilde{w}_{k-1})$ (11)
$=\tilde{w}_{k-1}-\eta{\mathcal{L}}^{\prime}(\tilde{w}_{k-1})+\eta\frac{w_{0}-w_{k}}{\eta}$
$=w_{k}+w_{0}-w_{k}$